# Hybrid Hidden Markov Model for Modeling Equity Excess Growth Rate Dynamics: A Discrete-State Approach with Jump-Diffusion

### 10 Mar 2026 v1

@@ 2.4. Factor models and multi-asset extensions
Modeling the joint dynamics of large asset universes requires a framework for managing cross-sectional dependence. The Single-Index Model (SIM) of Sharpe (Sharpe, 1963) provides the canonical linear factor decomposition: each asset’s excess return is decomposed into a common market factor component and an idiosyncratic residual, drastically reducing the number of parameters needed compared with a full multivariate model. The SIM remains foundational in portfolio construction and risk management despite the availability of richer multi-factor models such as Fama-French (Fama and French, 1993), because its simplicity enables transparent attribution and scalable computation. Existing HMM-based work in finance has been predominantly univariate, focusing on single-asset or single-index regime identification. Applications to individual equities (Nguyen, 2018), volatility modeling (Rossi and Gallo, 2006), and macro time series (Hamilton, 1989; Kim and Nelson, 1999) collectively demonstrate the utility of latent state representations but leave open the question of how HMM-derived generative models can be extended to large cross-sections of assets in a computationally tractable way. Mixture hidden Markov models (Dias et al., 2010) offer a partial solution by allowing multiple assets to share a common latent state structure, but estimating such models jointly across hundreds of assets is computationally demanding. An alternative approach to multi-asset dependence modeling uses copulas, which separate the specification of marginal distributions from the dependence structure (Embrechts et al., 2002; Cherubini et al., 2004). Copula-based generators can capture nonlinear and tail dependence that linear factor models miss, and they have been applied extensively in credit risk and portfolio simulation. However, copula estimation becomes computationally demanding in high dimensions, and the choice of copula family introduces additional model selection complexity. The present study addresses the multi-asset challenge through a complementary route by combining the univariate HMM framework with the SIM factor decomposition: the hybrid HMM is estimated on the SPY index, which serves as the market-factor generator, and asset-level paths are reconstructed via the linear SIM projection. This construction exploits the fact that most of the shared volatility clustering across equities is driven by the broad market, yielding a scalable generative pipeline for a 424-asset universe that inherits the stylized-facts fidelity of the HMM without requiring joint estimation of a high-dimensional model.
@@ 3. Methods
We computed the excess growth rate $G_{i,j}$ for ticker $i$ between time period $j-1\rightarrow{j}$ for a given equity price series $P_{i,j}$ (units: dollars per share (USD/share)). To measure how much each stock’s growth exceeded a risk-free baseline, we defined the excess growth rate as:
@@
where $\Delta t$ represented the time step set to $1/252$ for daily data (units: years), and $r_{f}$ denoted the constant continuously compounded risk-free rate (units: year-1) derived from Separate Trading of Registered Interest and Principal of Securities (STRIPS) bond yields observed on September 10, 2025. Unlike a simple return, the excess growth rate has units of inverse time (year-1) and directly quantifies the premium earned above a continuously compounded risk-free benchmark; log growth rates are also time-additive, making them consistent with a continuous compounding framework. While $r_{f}$ varied dynamically over the sample period, we used a constant proxy to isolate the volatility patterns in the data without adding noise from interest rate movements.
Discrete hidden Markov model.
@@
We defined the hidden Markov model using the tuple $\mathcal{M}=(\mathcal{S},\mathcal{O},\mathbf{T},\mathbf{E},\bar{\pi})$ (Rabiner and Juang, 1986). The hidden state $S_{t}\in\mathcal{S}=\{1,\dots,N\}$ represented market mood regimes and was defined by quantile bins of a fitted cumulative distribution function (CDF) for the excess growth rate series. The observation space $\mathcal{O}\subset\mathbb{R}$ consisted of continuous excess growth rate values, the transition matrix $\mathbf{T}$ governed regime-to-regime dynamics, the emission model $\mathbf{E}$ was parameterized by state-conditional Student-t distributions, and $\bar{\pi}$ denoted the stationary distribution. The state-conditional observation model specified that the excess growth rate, given hidden state $k$, followed a location-scale Student-t distribution with $\nu=5$ degrees of freedom:
(2) $G_{t}\mid S_{t}=k\sim\mu_{k}+\sigma_{k}\cdot t_{\nu},\qquad\nu=5.$
To map observations to hidden states during estimation, we defined quantile boundaries $Q_{k}=F_{L}^{-1}(k/N;\mu_{L},b_{L})$ for $k\in\{1,\dots,N-1\}$, with $Q_{0}=$-$\infty$ and $Q_{N}=+\infty$ so that the partition covers the full support of the distribution. An observation was assigned to state $k$ when $Q_{k-1}<G_{t}\leq Q_{k}$.
The Laplace distribution was chosen for state partitioning rather than for emission modeling. Its sharper peak at the mean better matched the concentration of small price movements observed in the SPY dataset (Mandelbrot, 1963; Toth and Jones, 2019; Cont, 2001; Kotz et al., 2001), and its closed-form quantile function supported the 424-ticker pipeline without iterative numerical optimization (Bilmes, 1998). The quantile boundaries were derived from a parametric Laplace fit and therefore did not enforce equal counts per bin in the historical data; instead, they provided a partition of the distribution that preserved tail shape. By using the Laplace distribution to define the bins, the model established a baseline characterization of the return distribution’s shape before the transition dynamics and jump mechanism were layered on top (Kou, 2002). The emission distribution within each state, however, was a location-scale Student-t with $\nu=5$ degrees of freedom, chosen because its heavier tails relative to the Normal distribution better reproduced the observed excess kurtosis of the growth rate series (a sensitivity analysis over $\nu\in\{3,4,5,6,7,8,10,15,30,\infty\}$ identified $\nu=5$ as the value that minimized the kurtosis gap without degrading distributional fidelity; $\nu=\infty$ recovers the Normal baseline).
@@
We estimated the transition matrix $\mathbf{T}$ through direct frequentist counting rather than iterative optimization. This approach differed from the traditional expectation-maximization (EM) algorithm used for parameter estimation in Gaussian mixture and hidden Markov models (Bilmes, 1998). However, the discrete state assignments allowed for straightforward counting of observed transitions between regimes, making the approach both conceptually simple and free of initialization sensitivity. A consequence of direct counting was that empirical transition matrices often had low probabilities for exiting central states into tail states, which caused models to lose the effect of volatility clustering too quickly (Bulla and Bulla, 2006). We corrected this by adding jump parameters $\epsilon$, representing the probability of a jump event, and $\lambda$, representing the mean duration, along with a tail-state selection rule. Tail-states were defined as $\mathcal{S}_{bottom}=\{1,\dots,N_{tail}\}$ and $\mathcal{S}_{top}=\{N-N_{tail}+1,\dots,N\}$, where $N_{tail}$ was user-configurable. If a jump was triggered, the jump-duration $K$ was sampled from a Poisson distribution such that $K\sim\text{Poisson}(\lambda)$, which was standard for modeling independent events in fixed intervals (Glasserman, 2003). During these $K$ steps, the standard Markovian transitions were overridden to target tail-states, with a user-configurable bias $p_{neg}$ (default 0.52) that favored negative-tail states to reproduce gain/loss asymmetry.
@@
To characterize the excess growth rates within each regime, we parameterized the emission model with location-scale Student-t distributions $\mu_{k}+\sigma_{k}\cdot t_{5}$, where $\mu_{k}$ and $\sigma_{k}$ were the sample mean and standard deviation of observations assigned to state $k$. During simulation, when the model was in state $k$, the continuous excess growth rate ($\hat{G}_{t}$) was sampled from the corresponding local distribution. This hybrid approach ensured that the model captured sudden large price moves while keeping smooth variation within each regime (Kou, 2002). The fitted Laplace CDF closely matched the empirical CDF for SPY, and the empirical transition matrix exhibited a dominant near-diagonal band reflecting short-range regime persistence (Figure S1, Online Appendix S1). Critically, under pure Markovian dynamics the model exited extreme (tail) states after only 1–2 steps on average, far shorter than the multi-week volatile episodes observed in real markets. This rapid reversion was the core limitation that motivated the jump-duration extension.
Growth rate generation.
The basic sampling procedure for the HMM (without jumps) proceeded as follows. The stationary distribution $\bar{\pi}$, representing the long-run probability of occupying each regime, satisfies the balance equation $\bar{\pi}=\bar{\pi}\mathbf{T}$ and was estimated numerically by raising the transition matrix to a large power ($\mathbf{T}^{50}$) (Hamilton, 2018; Grinstead and Snell, 1997). An initial hidden state was drawn from $X_{0}\sim\bar{\pi}$, then for each subsequent step $n=1,\dots,T$: (i) the next hidden state was sampled from the transition row $X_{n}\sim\mathbf{T}_{X_{n-1},\,\cdot}$, and (ii) an observation was sampled from the emission distribution $O_{n}\sim\mathbf{E}_{X_{n}}$. This produced a sequence of hidden states $\{X_{1},\dots,X_{T}\}$ and corresponding observations $\{O_{1},\dots,O_{T}\}$.
@@
Statistical validation metrics
To assess whether synthetic paths were statistically consistent with empirical data, we applied two nonparametric goodness-of-fit tests to the marginal distributions of simulated excess growth rate sequences. For each of the 1,000 simulated paths, the empirical cumulative distribution function (CDF) of the simulated sequence was compared against the empirical CDF of the historical in-sample or out-of-sample observations.
The Kolmogorov-Smirnov (KS) statistic is defined as:
(4) $D_{KS}=\sup_{x}\left\lvert F_{n}(x)-F_{m}(x)\right\rvert$
where $F_{n}$ and $F_{m}$ denote the empirical CDFs of the observed and simulated sequences of lengths $n$ and $m$, respectively (Kolmogorov, 1933; Smirnov, 1948). The null hypothesis of distributional equivalence is rejected at significance level $\alpha$ when $D_{KS}$ exceeds the critical value $c(\alpha)\sqrt{(n+m)/(nm)}$.
The Anderson-Darling (AD) statistic places greater weight on the distributional tails than the KS test. The original one-sample form (Anderson and Darling, 1952) measures deviation from a fully specified reference CDF $F_{0}$:
(5) $A^{2}=-n-\sum_{i=1}^{n}\frac{2i-1}{n}\left[\ln F_{0}(X_{(i)})+\ln\left(1-F_{0}(X_{(n+1-i)})\right)\right]$
where $X_{(1)}\leq\cdots\leq X_{(n)}$ are the order statistics of the sample. In this study, we applied the two-sample Anderson-Darling test, which replaces $F_{0}$ with the empirical CDF of the observed SPY series; critical values are obtained from the $k$-sample variant of Anderson and Darling as implemented in the HypothesisTests.jl library. The AD test is particularly suited for assessing tail-distributional fidelity, which is the primary concern for risk management applications.
We reported the pass rate as the proportion of the 1,000 simulated paths for which the null hypothesis of distributional equivalence was not rejected at the $\alpha=0.05$ significance level. A pass rate above 95 percent indicated that the synthetic paths were, in aggregate, statistically indistinguishable from the historical data under that test. Pass rates were reported separately for the in-sample and out-of-sample windows and for both model variants (HMM without jumps and the hybrid HMM with jumps). Both KS and AD tests assume i.i.d. observations; because volatility clustering induces positive autocorrelation in the tail-indicator process $\mathbf{1}(G_{t}\leq x)$, the effective sample size is smaller than $T$ and reported p-values may be slightly miscalibrated. The continuous distance metrics below are unaffected by this assumption.
Continuous distributional distance metrics
Binary pass rates summarize distributional fidelity but discard information about the magnitude of any discrepancy. The Wasserstein-1 distance ($W_{1}$) and the Hellinger distance ($H$) were computed alongside KS and AD to provide continuous, effect-size measures that do not depend on sample size.
The Wasserstein-1 distance measures the minimum work required to transform one distribution into the other, where work is distance times probability mass moved. For two equal-length empirical samples with order statistics $x_{(1)}\leq\cdots\leq x_{(T)}$ and $y_{(1)}\leq\cdots\leq y_{(T)}$, the one-dimensional Wasserstein-1 distance is:
(6) $W_{1}=\frac{1}{T}\sum_{i=1}^{T}\left\lvert x_{(i)}-y_{(i)}\right\rvert$
$W_{1}$ is expressed in the same units as the data (daily excess growth rates) and equals zero only when the empirical distributions are identical. Because it is computed from sorted order statistics, the temporal ordering of observations is irrelevant; smaller values indicate closer distributional agreement across the full quantile range.
The Hellinger distance measures distributional overlap and is bounded in $[0,1]$, with $H=0$ indicating identical distributions and $H=1$ indicating completely disjoint support. Given normalized histogram estimates $\{p_{k}\}_{k=1}^{K}$ and $\{q_{k}\}_{k=1}^{K}$ of the observed and simulated marginal densities over a common $K$-bin grid:
(7) $H(P,Q)=\frac{1}{\sqrt{2}}\sqrt{\sum_{k=1}^{K}\left(\sqrt{p_{k}}-\sqrt{q_{k}}\right)^{2}}$
The bounded scale makes $H$ directly comparable across models and evaluation windows without rescaling. Histograms were constructed on the common support of observed and simulated samples using $K=50$ equal-width bins. Both $W_{1}$ and $H$ were computed path-by-path and averaged across the 1,000 paths; standard errors were estimated as $\mathrm{std}/\sqrt{n_{\mathrm{paths}}}$.

### 02 Apr 2026 v2

@@ 2.4. Factor models and multi-asset extensions
Existing HMM-based work in finance has been predominantly univariate (Hamilton, 1989; Rossi and Gallo, 2006; Nguyen, 2018; Kim and Nelson, 1999), leaving open the question of how HMM-derived generative models can scale to large cross-sections of assets. One strategy could be to use an HMM-based approach to capture the marginal distribution of a market factor (e.g., SPY) and then reconstruct individual asset paths via a linear factor model. The Single-Index Model (SIM) of Sharpe (Sharpe, 1963) decomposes each asset’s excess return into a common market factor and an idiosyncratic residual, reducing the parameter count relative to a full multivariate model. Mixture hidden Markov models (Dias et al., 2009) allow multiple assets to share a latent state structure, but joint estimation across hundreds of assets is computationally demanding. On the other hand, copulas offer an alternative by separating marginal distributions from the dependence structure (Embrechts et al., 2002; Cherubini et al., 2004), though estimation grows expensive in high dimensions and the choice of copula family adds model selection complexity. The present study addresses the multi-asset challenge through both approaches: a SIM factor decomposition provides a scalable baseline for a 424-asset universe (the hybrid HMM generates market-factor paths from SPY, and asset-level paths are reconstructed via the linear projection), while copula-based dependence models (Gaussian, Student-t, and C-vine (Aas et al., 2009)) preserve each asset’s independently fitted HMM marginal and inject cross-asset dependence via rank reordering. The empirical comparison reveals that copulas substantially improve per-asset distributional accuracy relative to the SIM, and that a single Student-t copula is a parsimonious and sufficient specification for the equity pairs examined.
@@ 3. Methods
We computed the excess growth rate $G_{i,j}$ for ticker $i$ between time period $j-1\rightarrow{j}$ given the equity price series $P_{i,j}$:
@@
where $\Delta t=1/252$ (one trading day in years) and $r_{f}$ was a constant continuously compounded risk-free rate derived from STRIPS bond yields. The excess growth rate has units of year-1 and is time-additive under continuous compounding. A constant $r_{f}$ was used to isolate volatility patterns without adding noise from interest rate movements.
Discrete hidden Markov model growth rate generator.
@@
We defined the hidden Markov model as the tuple $\mathcal{M}=(\mathcal{S},\mathcal{O},\mathbf{T},\mathbf{E},\bar{\pi})$ (Rabiner and Juang, 1986), where the hidden state $S_{t}\in\mathcal{S}=\{1,\dots,N\}$ represented market regimes, the observation space $\mathcal{O}\subset\mathbb{R}$ consisted of continuous excess growth rate values, $\mathbf{T}$ governed regime-to-regime transitions, $\mathbf{E}$ specified the state-conditional emission model, and $\bar{\pi}$ denoted the stationary distribution. States were defined by quantile boundaries $Q_{k}=F_{L}^{-1}(k/N;\mu_{L},b_{L})$ of a fitted Laplace distribution, with $Q_{0}=$-$\infty$ and $Q_{N}=+\infty$; the Laplace was chosen for its sharper peak (better matching the concentration of small price movements (Mandelbrot, 1963; Cont, 2001; Kotz et al., 2001)) and its closed-form quantile function, which scaled to the 424-ticker pipeline without iterative optimization (Bilmes, 1998). An observation was assigned to state $k$ when $Q_{k-1}<G_{t}\leq Q_{k}$. The within-state emission followed a location-scale Student-t distribution:
(2) $G_{t}\mid S_{t}=k\sim\mu_{k}+\sigma_{k}\cdot t_{\nu},\qquad\nu=5,$
where $\nu=5$ was selected from a sensitivity analysis over $\nu\in\{3,4,5,6,7,8,10,15,30,\infty\}$ as the value that minimized the kurtosis gap without degrading distributional fidelity ($\nu=\infty$ recovers the Normal baseline).
@@
We estimated the transition matrix $\mathbf{T}$ through direct frequentist counting: the discrete state assignments allowed straightforward counting of observed transitions between regimes, making the approach conceptually simple and free of the initialization sensitivity inherent to EM (Bilmes, 1998). A consequence of direct counting was that empirical transition matrices often had low probabilities for exiting central states into tail states, which caused models to lose the effect of volatility clustering too quickly (Bulla and Bulla, 2006). We corrected this by adding jump parameters $\epsilon$, representing the probability of a jump event, and $\lambda$, representing the mean duration, along with a tail-state selection rule. Tail-states were defined as $\mathcal{S}_{bottom}=\{1,\dots,N_{tail}\}$ and $\mathcal{S}_{top}=\{N-N_{tail}+1,\dots,N\}$, where $N_{tail}$ was user-configurable. If a jump was triggered, the jump-duration $K$ was sampled from a Poisson distribution such that $K\sim\text{Poisson}(\lambda)$, which is standard for modeling independent events in fixed intervals (Glasserman, 2003). During these $K$ steps, the standard Markovian transitions were overridden to target tail-states, with a user-configurable bias $p_{neg}$ (default 0.52) that favored negative-tail states to reproduce gain/loss asymmetry (Figure 2).
@@
Given the state assignments and transition matrix, the remaining model components were estimated directly from the data. The emission parameters $\mu_{k}$ and $\sigma_{k}$ were the sample mean and standard deviation of observations assigned to state $k$. The empirical transition matrix exhibited a dominant near-diagonal band reflecting short-range regime persistence (Figure S1, Online Appendix S1). The stationary distribution $\bar{\pi}$ satisfied $\bar{\pi}=\bar{\pi}\mathbf{T}$ and was estimated by matrix powering ($\mathbf{T}^{50}$) (Hamilton, 2018; Grinstead and Snell, 1997). An initial state was drawn from $X_{0}\sim\bar{\pi}$; at each subsequent step the chain either transitioned according to the empirical row $\mathbf{T}_{X_{n-1},\,\cdot}$ (probability $1-\epsilon$) or entered a Poisson jump episode (probability $\epsilon$) that forced the state into the tail set $\mathcal{S}_{-}$ or $\mathcal{S}_{+}$ for $K\sim\text{Poisson}(\lambda)$ consecutive steps before normal transitions resumed. In both cases, an observation was sampled from the emission distribution $O_{n}\sim\mathbf{E}_{X_{n}}$. The full simulation algorithm and remaining pipeline stages (model construction, state decoding) are given in Online Appendix S4 (Algorithm 1).
@@
Validation metrics
Each of the 1,000 simulated paths was compared against the historical excess growth rate series using four complementary metrics. The two-sample Kolmogorov-Smirnov (KS) test (Kolmogorov, 1933; Smirnov, 1948) assessed overall distributional agreement; the two-sample Anderson-Darling (AD) test (Anderson and Darling, 1952) placed greater weight on the tails. For each test, we reported the pass rate: the proportion of paths for which the null hypothesis of distributional equivalence was not rejected at $\alpha=0.05$. To complement these binary verdicts with continuous effect-size measures, we also computed the Wasserstein-1 distance ($W_{1}$) and the Hellinger distance ($H\in[0,1]$), averaged path-by-path across simulations. Formal definitions of all four metrics are given in Supplemental Section S10.